Assume the current interest rate on a one-year Treasury bond ( ) is 1.10 percent, the current rate on a two-year Treasury bond (R2) is 1.26 percent, and the current rate on a three-year Treasury bond (1R3) is 1.37 percent. If the unbiased expectations theory of the term structure of interest rates is correct, what is the one-year interest rate expected on T-bills during year 3 (E3or 3)? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16)) Interest rate expected during year 3
Answer #1

3 year rate = 1.37%

2 year rate = 1.26%

Let the 1 year rate during year 3 is r%

(1 + 1.37%)3 = (1 +1.26%)2 * (1 + r%)

1.04166 = 1.025359 * (1 + r%)

1 + r% = 1.015904

r% = 1.59%